OpenStudy (anonymous):

HELP BECOMING FAN AND MEDALS PLZ 1.) Find the derivative of f(x) = -3/x at x = -4 a) 3/4 b) 3/16 c) 16/3 d) 4/3 2.) Find the limit of the function by using direct substitution. lim x-->0 (x^2-2) a) 2 b) -2 c) doesnt exist d)0

OpenStudy (anonymous):

@Miss.SweetiePie @ganeshie8

OpenStudy (anonymous):

@stratemeyerjr

OpenStudy (welshfella):

1. convert to negative power 3/x = 3x^-1 now to find the derivative use if y = ax^n derivative = anx^(n-1)

OpenStudy (astrophysics):

To make it more clear, \[\frac{ d }{ dx } x^{n} = nx^{n-1}\] this is the power rule apply it, find the derivative using this then plug in x = -4.

OpenStudy (welshfella):

here a = 3 and n = -1

OpenStudy (astrophysics):

And for 2, do as it says :)

OpenStudy (astrophysics):

\[\lim_{x \rightarrow 0} ~~ x^2-2\] implying x is going to 0, so that should be an obvious substitution.

OpenStudy (anonymous):

for 1 i got 3/16 is that right?? @welshfella and number 2 i got B -2

OpenStudy (welshfella):

1. derivative = 2 / x^2 when x = -4 deriavative = 3 / 16 yes correct

OpenStudy (welshfella):

also b = (0)^2 - 2 = -2 yes

OpenStudy (welshfella):

good job

OpenStudy (anonymous):

thank u @welshfella

OpenStudy (welshfella):

yw